OpenStudy (anonymous):

Find the roots of the polynomial equation. x2 + 8x – 48 = 0

OpenStudy (linn99123):

x² - 8x - 48 = 0 x² - 4x = 48 + 4² x² - 4x = 48 + 16 (x - 4)² = 64 x - 4 = 8 x = 8 + 4, x = 12 x = - 8 + 4, x = - 4 Answer: b) 12 and - 4

OpenStudy (anonymous):

thats just quadratic. you can either factor, complete the square, or use quadratic formula on quadratics.

OpenStudy (anonymous):

how did you know the answer choices

OpenStudy (anonymous):

because he completed to square and solved it.